If 𝐴 = 3π‘₯ 2 + 6𝑦 𝑖 βˆ’ 14𝑦𝑧𝑗 + 20π‘₯𝑧 2π‘˜, οΏ½οΏ½Χ¬ οΏ½οΏ½evaluate 𝐴 βˆ™ π‘‘π‘Ÿ from (0,0,0) to (1,1,1) along the following path C: π‘Ž. π‘₯ = 𝑑, 𝑦 = 𝑑 2 , 𝑧 = 𝑑 3 b. the straight lines from (0,0,0) to (1,0,0) and then to (1,1,0) and the to (1, 1, 1) c. the straight line joining (0,0,0) to (1,1,1).


1
Expert's answer
2021-05-10T18:32:38-0400

(a) Give, "x=t,y=t^2,z=t^3"


The points corresponds to t=1


"dx=dt,dy=2tdt,dz=3t^2dt


\\[9pt]
\int \vec{A}.\vec{dr}=\int(3x^2+6y)dx-14yzdy+20xz^2dz
\\[9pt]


 =\int_0^1(3t^2+6t^2)dt-14t^5(2tdt)+20t^7(3t^2dt)
\\[9pt]
 =5"


(b)

"\int_{(0,0,0,)}^{(1,0,0)} \vec{A}.\vec{dr}=\int_0^1(3x^2+6y)dx-\int_0^014yzdy+\int_0^020xz^2dz
 \\[9pt]=x^3+6xy|_0^1=1"


"\int_{(1,1,0,)}^{(1,1,1)} \vec{A}.\vec{dr}=\int_1^1(3x^2+6y)dx-\int_1^114yzdy+\int_0^120xz^2dz
 \\[9pt]=20\dfrac{xz^3}{3}|_0^1=0"


(c)


"\int_{(0,0,0,)}^{(1,1,1)} \vec{A}.\vec{dr}=\int_0^1(3x^2+6y)dx-\int_0^114yzdy+\int_0^120xz^2dz

 \\[9pt]=x^3+6xy|_0^1-7y^2z|_0^1+\dfrac{20xz^3}{3}|_0^1\\[9pt]=1+6-7+\dfrac{20}{3}=\dfrac{20}{3}"